OpenStudy (anonymous):

Compare and contrast the two quadratic equations below. In order to receive full credit, use complete sentences to describe the following: The direction each parabola opens The vertex of each parabola y = x2 − 2x y = −2x2 + 4x − 3

OpenStudy (anonymous):

What is your mean ??? In y = x2 − 2x we have : x2-2x=0 y=0 In y = −2x2 + 4x − 3 we have : -4x+4x-3 = -3 y=-3